Questions E and F please

JUST DEW IT CORPORATION 2017 and 2018 Balance Sheets Assets Liabilities and Owners Equity 2017 2018 2017 2018 Current assets Current liabilities Cash Accounts $ 13,280 14,600 19,200 51,000 $63,200 84,800 Accounts payable Notes payable $ 40,000 $41,800 21,600 7,520 14,240 receivable Inventory 42,400 Total Total $54,240 63,400 Long-term debt $32,000 24,000 Owners equity Common stock and paid-in $ 40,000 40,000 surplus 272,600 $233,760 312,600 Total liabilities and owners $320,000 $400,000 Retained earnings 193,760 Net plant and equipment $256,800 315,200 Total Total assets $320,000 $400,000 equity Based on the balance sheets given for Just Dew It a. Calculate the current ratio for each year. (Do not round intermediate calculations b. Calculate the quick ratio for each year. (Do not round intermediate calculations and c. Calculate the cash ratio for each year. (Do not round intermediate calculations and d. Calculate the NWC to total assets ratio for each year. (Do not round intermediate e. Calculate the debt-equity ratio and equity multiplier for each year. (Do not round f. Calculate the total debt ratio and long-term debt ratio for each year. (Do not round and round your answers to 2 decimal places, e.g., 32.16.) round your answers to 2 decimal places, e.g.,32.16.) round your answers to 2 decimal places, e.g., 32.16.) calculations. Enter your answers as a percent rounded to 2 decimal places, e.g 32.16.) intermediate calculations and round your answers to 2 decimal places, e.g., 32.16 intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)
Answer #1

e.

Debt-equity ratio = Total Liabilities/Total equity

Debt-Equity Ratio for 2017 = (54240+32000)/233760

Debt-Equity Ratio for 2017 =0.37

Debt-Equity Ratio for 2018 = (63400+24000)/312600

Debt-Equity Ratio for 2018 = 0.28

Equity Multiplier = Total Assets/Stockholder's Equity

Equity Multiplier for 2017 = 320000/233760

Equity Multiplier for 2017 = 1.37

Equity Multiplier for 2018 = 400000/312600

Equity Multiplier for 2018 = 1.28

f.

Total Debt Ratio = Total Debt/Total Assets

Debt Ratio for 2017 = (54240+32000)/320000

Debt Ratio for 2017 = 0.27

Debt Ratio for 2018 = (63400+24000)/400000

Debt ratio for 2018 = 0.22

Long Term Debt Ratio = Long Term Debt/Total Assets

Long Term Debt Ratio for 2017 = 32000/320000

Long Term Debt Ratio for 2017 = 0.10

Long Term Debt Ratio for 2018 = 24000/400000

Long Term Debt Ratio for 2018 = 0.06